OpenStudy (anonymous):

Factor the trinomial completely.. x^2^n+8x^n+15

OpenStudy (anonymous):

\[ a^2 + 8 a +15=(a+5)(a+3) \]

OpenStudy (anonymous):

\[ (x^n)^{2}+8x^n+15= (x^n+3)(x^n+5) \]